OpenStudy (anonymous):

The difference of two positive numbers is 3 and the sum of their squares is 89. Find the numbers.

OpenStudy (aroub):

The numbers are 5 and 8. 8 - 5 = 3 8^2 = 64 5^2 = 25 64 + 25 = 89

OpenStudy (anonymous):

thank you

OpenStudy (aroub):

You're welcome :)